A t-shirt store in a monopolistically competitive market faces a demand curve for t-shirts given by P = 25 – 0.5Q (and so a marginal revenue of MR = 25 – Q). The variable costs of producing a t-shirt are VC = 3Q and so the marginal costs are constant at $3. If the t-shirt store is in a long-run equilibrium, what must its fixed costs be?

$308

$242

$66

$418

$14
Answer #1

$242 ----- is correct

Profit is maximized at MR = MC

25 - Q = 3

Q= 22

P= 25 - 0.5*22= 14

Profit is 0 in long run equilibrium

Profit = PQ - VC - FC

0 = 14*22 - 3*22 - FC

FC = 242